20190227 DX Night INAF Science

01:35 Josh start pointing and collimation work

02:04 OSCO complete successfully, with the exception the GOPT did not start. First object is 15 mag so continueing.
  • I created the c2m matrix via AO> help, (ee->control())->c2m()
  • The Optical diag. for Gopt still is not running properly. No data set saved (no save data command on AdSec Control)
  • Will look into this.

02:19 Point and collimate near first science object

02:38 AO Loop Closed
  • Bin 4
  • Optimize gain gave 3.7, 3.7, 3.7
    • Loop opened with skip frame. Close loop from FLAO GUI
    • Set Manually to 2.0, 1.5, 1.5
  • TN 20190228_023949

11:40 CCD39 when "Not Live"
  • Died on PresetAO
  • Error Reading Shared buffer masterdiagnostic...(-5001) Timeout_Error
    • Stopped CCD39 Controler
    • Stopped Slope Computer Controller
    • Cycled power on Little Joe 39
    • Started slope Computer controller (Init state)
    • Started CCD39 Computer Controllor (slope initialized)
    • Set Binning to 1x1: success
    • Resent Preset from LUCI2: Success

-- DougMiller - 28 Feb 2019
Topic revision: r2 - 28 Feb 2019, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback